# 3D Beamforming Through Joint Phase-Time Arrays

### v1

@@ IV-B Iterative Results
Within this segment, we present a performance evaluation of the algorithms introduced in Secs. III-C and III-D. Fig. 6 provides a comparison between the greedy and gradient descent algorithms with joint and separated optimization scenarios. It’s worth noting that the disparities in beamforming gains between separated and joint optimizations exhibit a consistent pattern, as discussed previously.
@@
In Fig. 4(a), we present the average gains for each user when their bandwidth allocation is uniform. Through joint optimization, both the greedy and gradient descent algorithms achieve an average gain increase of $2.26$ dB and $2.11$ dB, respectively. A similar trend is observed when we consider different bandwidth allocations in Fig. 4(b). It’s worth noting that the greedy and gradient descent algorithms yield comparable performance. The primary distinction between these two methods lies in their runtime, with the gradient descent algorithm running approximately 250x faster in joint optimization case.
@@
In the realm of 2D frequency-dependent beam design, the algorithm proposed by [5] stands as the current state-of-the-art. We have extended their iterative algorithm to 3D for the sake of comparative analysis. The results of this comparison are depicted in Fig. 6 as JPTA Iterative, which are across various bandwidth allocations for a scenario with $N_{u}=2$. It’s important to highlight that when the disparity in bandwidth allocation is relatively small ($\alpha>0.35$), all algorithms employing the joint optimization scheme tend to converge to the same outcome. This observation also underscores a fundamental limitation of the separated optimization approach.
@@
Moreover, our one-shot analytical solution surpasses the iterative state-of-the-art approach proposed by [5] for scenarios with unequal bandwidth allocations ($\alpha\leq 0.35$), achieving up to a $1.54$ dB gain. For $\alpha\leq 0.2$, even the use of separated optimization within the gradient descent algorithm outperforms the other methods, with the exception of the gradient descent algorithm using joint optimization. Additionally, the run-time of the gradient descent algorithm is at least 5 times faster than the state-of-the-art. In conclusion, the gradient descent algorithm with joint optimization consistently performs well across various bandwidth allocations, thus offering more reliable results while running faster than alternative iterative algorithms. When the bandwidth is divided fairly, we can achieve similar performance to the gradient descent algorithm using a one-shot joint analytical derivation with LS.
@@ V Conclusion
In this paper, we have investigated 3D frequency-dependent beam design to serve multiple users. We have provided analytical derivations to maximize the beamforming gain. Subsequently, we have enhanced our analytical results by using two iterative algorithms: greedy and gradient descent. Through simulations of 3D beamforming gain, we have analyzed the performance of our proposed algorithms in comparison to the state-of-the-art, which was originally proposed for 2D and extended to 3D by us.
Our findings indicate that our proposed gradient descent algorithm surpasses the state-of-the-art, delivering more reliable results across various scenarios. Furthermore, our joint analytical derivation matches the performance of the gradient descent algorithm when the bandwidth allocation is relatively uniform. The simulation results clearly show that the separated design, which is usually adopted for the UPA 3D beam design, does not work well for the JPTA use case.

### v3

@@ IV-B Results of iterative algorithms
This segment presents a performance evaluation of the algorithms introduced in Secs. III-C and III-D. Fig. 5 compares the greedy and gradient descent algorithms with joint and separated optimization scenarios. It’s worth noting that the disparities in beamforming gains between separated and joint optimizations exhibit a consistent pattern, as discussed previously.
@@
In Fig. 5(a), we present the average gains for each user when their BW allocation is uniform. Through joint optimization, both the greedy and gradient descent algorithms achieve an average gain increase of $2.26$ dB and $2.11$ dB, respectively. A similar trend is observed when considering different BW allocations in Fig. 5(b). It’s worth noting that the greedy and gradient descent algorithms yield comparable performance. The primary distinction between these two methods lies in their runtime, with the gradient descent algorithm running approximately 250x faster in joint optimization cases.
@@
In the realm of 2D frequency-dependent beam design, the algorithm proposed by [5] stands as the current state-of-the-art. We have extended their iterative algorithm to 3D for the sake of comparative analysis. The results of this comparison are depicted in Fig. 6 as JPTA Iterative, which are across various BW allocations for a scenario with $N_{u}=2$. It’s important to highlight that when the disparity in BW allocation is relatively small ($\alpha>0.35$), all algorithms employing the joint optimization scheme tend to converge to the same outcome. This observation also underscores a fundamental limitation of the separated optimization approach.
@@
Moreover, our one-shot analytical solution surpasses the iterative state-of-the-art approach proposed by [5] for scenarios with unequal BW allocations ($\alpha\leq 0.35$), achieving up to a $1.54$ dB gain. For $\alpha\leq 0.2$, even the gradient descent algorithm with separated optimization outperforms the other methods, except for the gradient descent algorithm using joint optimization. In conclusion, the gradient descent algorithm with joint optimization consistently performs well across various BW allocations, thus offering more reliable results. When the BW is divided fairly, it can achieve a performance similar to the gradient descent algorithm using a one-shot joint analytical derivation with LS.
Figure 7: Time complexity of the proposed gradient descent method and the state-of-the-art iterative algorithm for different BW allocation scenarios. The codes are run on a server with Platium 8268 CPU @2.9GHz, and 2TB memory.
IV-C Time Complexity Analysis
In Fig. 7, we demonstrate the time complexity of the iterative algorithm by [5] and our proposed gradient descent method with different BW allocations scenarios. It’s important to note that run time of both algorithms are higher in average with unequal BW allocation ($\alpha\leq 0.35$). Since gradient descent does not improve upon the LS for equal BW allocation scenarios (see Fig. 6), the run-time is negligible. Lastly, average run-time of the gradient descent ($1.05$ s) is at least 5 times faster than the iterative algorithm ($5.8$ s).
@@ V Conclusion
In this paper, we have investigated the 3D frequency-dependent beam design with joint phase-time arrays, aiming to serve multiple users in a single shot and thereby reduce beam search and scheduling latency. We have presented analytical derivations to maximize the beamforming gain and then bolstered our findings with two iterative algorithms: greedy and gradient descent.
By simulating the 3D beamforming gain, we have compared the performance of our proposed algorithms with the state-of-the-art, originally designed for 2D and extended to 3D by us. Our findings indicate that our proposed gradient descent algorithm surpasses the state-of-the-art, delivering more reliable results across various scenarios. Furthermore, our joint analytical derivation matches the performance of the gradient descent algorithm when the bandwidth allocation is relatively uniform. Lastly, the simulation results clearly show that the separated optimization, usually adopted for the UPA 3D beam design, does not work well for the JPTA use case, since it leads to lower gain and more sidelobes.
Besides the beam design, a big challenge for the actual implementation of 3D JPTA is the aperture size and cost of delay elements. The number of delay elements increases linearly with the number of antennas in both azimuth and elevation directions in the considered structure. A future direction is to assign fewer delay elements in the elevation direction by connecting multiple antennas to a single delay element. Because the users are more distributed in the azimuth direction in the outdoor cellular network, this simplified design can still achieve most of the benefits of 3D beamforming.